OpenStudy (anonymous):

A circle is growing so that each side is increasing at the rate of 5 cm/min. How fast is the area of the circle changing at the instant the radius is 20 cm? Include units in your answer. I know this is a related rates question but it seems odd is all sides refering to a linear scale? so then it would correlate with the radius, or the circumference and if circumference i am having trouble getting a dc/dt to work so i can sub in the 5cm/min. please help!

jhonyy9 (jhonyy9):

side of a circle = perimeter - yes ?

OpenStudy (anonymous):

Thats what im geussing its just the all sides is odd, probably some bad coding on the test makesr part

OpenStudy (anonymous):

The problem im having is finding a format that gets me to DC/DT so i can sub in the 5cm/min

OpenStudy (anonymous):

the closest i have gotten is something like\[a=\pi * (\frac{ c^2 }{ 2\pi^2 })\]

OpenStudy (anonymous):

but i think i may be wrong here

OpenStudy (anonymous):

im not sure how to go about it once i get da/dt = pi d/dt (c^2/2pi^2)

OpenStudy (anonymous):

Like seperating the fraction

OpenStudy (anonymous):

i need an equation with DC/DT R and DA/DT somehow

OpenStudy (anonymous):

i think, or im doing this all wrong would be nice to get some input

OpenStudy (mathmale):

This is a nice "related rates" problem. Please go back and check the wording of the original problem. You've written, "a circle is growing so that each side increases in length at the rate of 5 cm/min. But circles don't have sides, do they?

OpenStudy (anonymous):

thats a direct copy-paste

OpenStudy (anonymous):

it has to mean circumference

OpenStudy (anonymous):

anything else doesnt make sense

OpenStudy (mathmale):

If that's the case, it looks as tho' you have 2 choices: Replace "circle" with "square" or solve the problem "a circle's circumference is increasing at the rate of 5 cm per min, and ... "

OpenStudy (mathmale):

Can you draw a circle with four sides? I can't.

OpenStudy (anonymous):

im trying to solve with circumference but am having difficulty formulating a problem

OpenStudy (mathmale):

C = 2*Pi*r A = Pi*r^2 Find the derivative with respect to time of each of the above. Don't forget to apply the Chain Rule.

OpenStudy (anonymous):

A= 2 PI c/2pi^2?

OpenStudy (anonymous):

but then would how do i derive it to get d/c/dt?

OpenStudy (anonymous):

and if thats the case where does r go?

OpenStudy (anonymous):

ive literally been trying to figure this out for 2 hours

OpenStudy (anonymous):

dc/dt*

OpenStudy (mathmale):

We're getting there. Again, A=pi*r^2. Differentiating with respect to time, t, we get \[\frac{ dA }{ dt}=2\pi r \frac{ dr }{ dt }\]

OpenStudy (mathmale):

Note: this is an application of the chain rule.

OpenStudy (mathmale):

Now, consider C=circumference=2*pi*r

OpenStudy (anonymous):

Yes, I got that on paper just didnt post it, thought it was a dead end

OpenStudy (mathmale):

find\[\frac{ dC }{ dt }\]

OpenStudy (anonymous):

Ahh

OpenStudy (mathmale):

You'll need to use the Chain Rule here.

OpenStudy (anonymous):

deriving 2 pi r gives me 1 or zero depending on if i sub in the r

OpenStudy (anonymous):

can i just sub 2pi r for dcdt then?

OpenStudy (anonymous):

rather it would be dc/dt=2pi

OpenStudy (anonymous):

is that correct?

OpenStudy (anonymous):

Remember i do not even have a dr/dt unless i just asume 5cm/min is dr/dt, in which case this whole circumference buissness is poitnless

OpenStudy (anonymous):

or would it be dc/dt=2pi dr/dt

OpenStudy (mathmale):

Please let's go thru this problem again in summary fashion: "solve the problem "a circle's circumference is increasing at the rate of 5 cm per min. How fast is the area of the circle increasing when t=3 sec?" First, we need to relate circumference to area, since we want to know how fast the area is increasing at time t=3. Circumference: C=2pi*r Area: A=pi*r^2 Let's solve for A in terms of C. Are you following all right so far?

OpenStudy (anonymous):

ok ok ok i have you know

OpenStudy (mathmale):

What do we already know? First, the circumference, C, is increasing at the rate of 5 cm/min Look, I have explained my strong belief that the problem as originally stated does not make sense, because a circle does not have sides. I invented a related problem in the hope we could go thru it quickly and then apply the method to a correction of the original problem. Thus, t=3 was an invention. I have serious issues with your "where the hell ... " question and your demand that I go back and read your problem. If you'd prefer to work with another person, fine.

OpenStudy (mathmale):

I know you're frustrated, but I will not long tolerate your taking your frustration out on me. Please make up your mind: do you want to continue or do you want to take a break and then let someone else try to help you?

OpenStudy (anonymous):

When you say "let's go thru this problem again" and not "lets look at a similar example" its extremely confusing, im right there with you on the fact that it is incorrectly formatted, but that was equally as confusing as the problem itself

OpenStudy (anonymous):

notice the phrase "this problem"

OpenStudy (anonymous):

I would greatly apreciate the help and do apologize for my temperment but ive had a long day

zepdrix (zepdrix):

I really really think they meant radius by "side length". I seriously doubt they were meaning for that to be circumference. Poorly written problem unfortunately, so it's hard to say.

OpenStudy (mathmale):

Sorry, but I am not responsible if the original problem I posted makes no sense. If you have a flesh-and-blood teacher, I'd suggest you take that up with him / her. Also sorry, but with the tone of this conversation being what it is, I'd prefer not to continue working with you, for the time being, at least. zepdrix has chosen to become involved. He is very capable and should be able to help you. Thank you, @zepdrix.

zepdrix (zepdrix):

:3

OpenStudy (anonymous):

Okay so we are dealing with a square then

OpenStudy (anonymous):

that makes a whole lot more sense

zepdrix (zepdrix):

Probably a squarcle jk XD

OpenStudy (anonymous):

yeah lol, honestly for all the work i put in i have a 1/3 chance of choosing the right starting point and there are 6 total questions and i have a chapter exam next lol

OpenStudy (anonymous):

im ganna leave it with a note that the question makes little logical sense and move on, i have alot more ahead of me, so its not worth me wasting anymore of mine or especailly your time for a fruitless endevour

zepdrix (zepdrix):

If it was meant to be the rate of change of circumference, \(\large\rm C'=5\) Then we first need to get the C value, \(\large\rm C=2\pi r=40\pi\) And replace the r with something involving C, \(\large\rm A=\pi r^2=\pi\left(\frac{C}{2\pi}\right)^2\) Then differentiate and plug in the pieces, yes? :)

zepdrix (zepdrix):

fine fine fine :) Whatever you gotta do hehe

OpenStudy (anonymous):

i got that exactly, but then realised i was missing r

zepdrix (zepdrix):

If it was meant to be radial change, \(\large\rm r'=5\) then of course the problem works out a little bit easier,\[\large\rm A=\pi r^2\]\[\large\rm A'=2\pi r\cdot r'\]

OpenStudy (anonymous):

also how would you preform the d/dt for a fraction like that im a bit rusty

OpenStudy (anonymous):

yeah that one is easy butt then i get a huge number

OpenStudy (anonymous):

626 per min

OpenStudy (anonymous):

628*

zepdrix (zepdrix):

I don't know why you think having no r is problem. We replaced it with C. But anyway, notice the denominator is constant :) Nothing to worry about there, just pull it aside before differentiating,\[\large\rm A=\frac{\pi}{4\pi^2}C^2\]And apply simple power rule to the C.

OpenStudy (mathmale):

@memman, it's hard to address your "how do I perform" question without actually seeing your work. It may take a while to share your work, but save time in the long run.

zepdrix (zepdrix):

\[\large\rm A=\pi\left(\frac{C}{2\pi}\right)^2\quad=\quad\pi\left(\frac{C^2}{2^2\pi^2}\right)\quad=\quad \frac{\pi}{2^2\pi^2}C^2\quad=\quad number\cdot C^2\]Just power rule, yes? :)

zepdrix (zepdrix):

You can do it merman! Just dive deep into the ocean! Find the pearl that you need!

zepdrix (zepdrix):

Long day? Humor gone? Ok my bad.

OpenStudy (anonymous):

So then after deriving i get ac/dt=2c*dc/dt

OpenStudy (anonymous):

no i apreciate it anything to lighten the mood

zepdrix (zepdrix):

I'm not sure what ac/dt is. Is that supposed to be our area derivative, dA/dt?

zepdrix (zepdrix):

I like the sneaky shortcut derivative notation myself, those primes are so clean. Whatever works for you though. Probably good to get comfortable with Leibniz notation for these types of problems.

OpenStudy (anonymous):

whoops da/dt

OpenStudy (anonymous):

yeah MyB

zepdrix (zepdrix):

\[\large\rm A=\frac{\pi}{2^2\pi^2}C^2\]So then,\[\large\rm A'=\frac{\pi}{2^2\pi^2}\cdot2C\cdot C'\]And earlier we determined that when the radius is 20, C=20 and C'=5 plug and chug

zepdrix (zepdrix):

woops, C=40pi*

zepdrix (zepdrix):

So ya, your derivative looks good, just don't forget about the junk in front.

zepdrix (zepdrix):

But again, I really really think they meant for r'=5 XD

zepdrix (zepdrix):

When I think of something growing in every direction, the "sides" growing, I'm thinking of it growing radially, outward. It certainly could be interpreted as circumference though :)

OpenStudy (anonymous):

ok well atleast i know now, i already submitted it like 5 min ago but i needed closure

zepdrix (zepdrix):

I guess you would want to compare it to similar problems that you've been working on. It should match the level of difficulty of those problems.

OpenStudy (anonymous):

Exactly its still useful material you guys are great, i also havent eaten since late lunch/dinner yesterday

OpenStudy (anonymous):

i got a little involved in this....

OpenStudy (anonymous):

been going since like 8 am

zepdrix (zepdrix):

oo that's no good. swim around, go find an eel or something D: gotta stay fed

zepdrix (zepdrix):

yikes

OpenStudy (anonymous):

you deserve a million medals my freind

OpenStudy (anonymous):

both of you

zepdrix (zepdrix):

np :D Study, do well, be blessed.

zepdrix (zepdrix):

LOL that picture XD